# A Survey on Inference Optimization Techniques for Mixture of Experts Models

### v1

@@ 3.2.2. Expert Quantization
In addition to model pruning, quantization is an effective method for reducing model size by converting high-precision weights into low-precision ones. Due to the redundancy of experts in MoE models, current studies focus on quantizing the weights of MoE experts. As shown in Figure 4-(b), experts are quantized into suitable low precision versions using different methods. MC-MoE (huang2024mc, ) uses both the frequency and scores of expert activations, along with the associated quantization loss, as metrics to determine the optimal quantization configuration for experts through a Linear Programming model. MoE-CSP (kim2022says, ) quantizes expert weights into 4 bits and designs specific cuda kernel to accelerate the computation. MoQE (kim2023mixture, ) quantizes expert weights to 2 bits to address the memory and latency challenges of MoE models, while QMoE (frantar2023qmoe, ) and CMoE (yuancompressed, ) compress MoE models to 1 bit, drastically reducing model size. Furthermore, some MoE-optimized systems leverage quantization for efficient system utilization. MoE-MPTQS (imani2024mixture, ) and HOBBIT (tang2024hobbit, ) dynamically select quantized experts to replace the original experts based on the current inputs, reducing the expert loading cost on memory-limited devices. EdgeMoE (yi2023edgemoe, ) determines the appropriate expert bit width by profiling expert importance on a calibration dataset statistically. Additionally, QMoE-Benchmark (li2024examining, ) provides a benchmark that explores several MoE structure-aware quantization heuristics, ranging from coarse to fine granularity. The study reveals that different MoE structures (e.g., blocks, experts, linear layers) require varying numbers of weight bits for effective and efficient quantization.
@@ 3.2.3. Expert Distillation
Knowledge distillation is an effective method for creating smaller, yet powerful models from larger ones. Due to the large size of MoE models, direct deployment is challenging. Therefore, As shown in Figure 4-(c), knowledge distillation offers a potential solution to generate a compact, high-performance model from the original MoE model. LLaVA-MoD (shu2024llava, ) leverages the MoE structure to balance scale reduction while maintaining the ability to capture complex multimodal knowledge through two consecutive distillation stages: mimic distillation and preference distillation. DeepSpeed-MoE (DeepSpeedmoe, ) employs staged knowledge distillation to create a distilled version of its proposed PR-MoE model, called MoS, which reduces model size while preserving performance. Furthermore, some studies focus on transferring the power of sparse models to dense models through knowledge distillation for more efficient deployment. For example, MoE-KD (salinas2022knowledge, ) and CMoE (yuancompressed, ) distill MoE-base speech recognition models into dense models, accelerating speech recognition process. OneS (xue2022one, ) merges experts into a single expert using various aggregation methods, and then distills the merged model with the original MoE model. Switch Transformers (fedus2022switch, ) and ELSM (artetxe2021efficient, ) also study the distillation techniques to distill their sparse model into a dense model. And Additionally, to simplify the construction of MoE models, LaDiMo (kim2024ladimo, ) converts a dense model into a sparse MoE model through layer-wise distillation, where each expert learns to approximate the results of the original layers.
@@ 3.2.4. Expert Decomposition
As shown in Figure 4-(d), Low-rank decomposition is another effective method for reducing model size by decomposing a large weight matrix into two smaller matrices. MPOE (gao2022parameter, ) employs the matrix product operator (MPO), a tensor decomposition technique derived from quantum many-body physics, to decompose expert weights into a central tensor and several auxiliary tensors. All experts in one layer share the same central tensor, thereby reducing the total number of parameters. MC-SMoE (li2023merge, ) applies low-rank decomposition to its merged experts, further reducing the model size. And MoE-$I^{2}$ (yang2024moe, ) identifies the importance of each expert and assigns higher ranks to more important experts while assigning lower ranks to less important ones for low-rank decomposition.
3.3. Algorithm Optimization
The core algorithm design of MoE is expert networks, and selecting appropriate granularity sparsity is a prerequisite for effectively utilizing the advantages of MoE. Next, we will fully explore the inference optimization of MoE models from three different strategies.
Figure 5. Algorithm improvement on experts.
@@ 3.3.1. Dynamic Gating
Given the significant progress in utilizing the sparsity of MoE models, an important question arises: how can we further leverage this sparsity during inference? Due to the vertical parallel structure of MoE experts, dynamically reducing the number of experts activated for each token clearly presents an effective strategy. Figure 5-(a) illustrates the MoE layer’s calculation process after the skip mechanism is applied, in contrast to the original process.
Li et al. (li2023adaptive, ) proposes a self-adaptive gating mechanism that dynamically determines the number of experts required for each token, based on the expert probability distribution. This method enhances training efficiency while preserving the sparsity of the MoE model and further reduces training time through curriculum learning. DynMoE (guo2024dynamic, ) introduces two innovative methods for expert activation: the top-down gating method, which allows different tokens to activate varying numbers of experts, and an adaptive process that dynamically adjusts the number of activated experts based on need. XMoE (yang2024xmoe, ) employs a strategy of using smaller, but more experts, dynamically selecting the number of activated experts based on a set threshold. AdapMoE (zhong2024adapmoe, ) is a co-design framework aimed at improving inference efficiency on edge devices. It adaptively applies three techniques: expert activation, expert prefetching, and GPU cache allocation. For deployment, MoE-Deploy (huang2023towards, ) integrates a dynamic gating method with the other two technologies to optimize performance.
3.3.2. Expert Merging
What other methods can we use to reduce the number of experts while maintaining or even improving their performance? One potential approach is inspired by ensemble learning, where merging outdated sub-experts based on their parameters could achieve this goal. Figure 5-(b) illustrates the expert merging process to decrease the number of parameters during inference.
Branch-Train-Merge (li2022branch, ) independently trains different parts of the model on distinct subsets of data, eliminating the need for large-scale multi-node synchronization typically required in traditional LLM training. Building on this, Branch-Train-Mix (sukhbaatar2024branch, ) merges the parameters of the MoE layer to create a unified model that can be further trained. Park et al. (park2024learning, ) observed that introducing a shared layer in the MoE could lead to performance degradation. To address this, they trained merged experts that learned the same features in different ways, improving their ability to generalize and mitigating catastrophic forgetting during incremental learning of multi-domain tasks. HC-SMoE (chen2024retraining, ) is a framework that uses hierarchical clustering to merge experts without requiring retraining, and can be applied in a task-agnostic manner. During inference, MEO (he2023merging, ) first merges the selected expert parameters and then performs efficient computation. FoE (wang2023fusing, ) fuses the outputs of expert models by leveraging their complementary knowledge of the data distribution, framing it as a supervised learning instance.
@@ 3.3.3. Sparse to Dense
In certain scenarios, dense models offer unique advantages due to their smaller number of parameters. So why not transform MoE models into dense target models all at once, achieving maximum sparsity while maintaining model performance? Most approaches use knowledge distillation to achieve this sparse-to-dense conversion, as illustrated in Figure 5-(c).
XFT (ding2024xft, ) proposes a novel method to supervise fine-tune dense LLMs. They first generate a sparse-upcycled MoE model, then transform it back into an efficient dense LLM of the same size and structure through a learnable merging mechanism. Switch Transformers (fedus2022switch, ) explores distillation techniques to convert a sparse model into a dense one, demonstrating that the dense model can retain over 30% of its performance even after compressing 97% of the parameters from the sparse MoE model. In another study ELSM (artetxe2021efficient, ), it also finds that a dense student model distilled from a sparse MoE model can improve performance. OneS (xue2022one, ) employs four distinct methods to generate the dense model, including summation, averaging, Top-K Knowledge Gathering, and Singular Value Decomposition Knowledge Gathering. The model is then refined to reduce noise by gathering sparse knowledge. TSEP (chen2022task, ) progressively eliminates non-expert components based on specific downstream tasks, ultimately converting the sparse MoE model into a dense counterpart. EWA (huang2023experts, ) replaces FFNs with specially designed MoEs during training, and then reverts to the original dense ViT for inference.
@@ 4. System-level Optimizations
Communication, wnode, l sep=5mm,calign=child edge, [Tutel (hwang2023tutel, )  Janus (10.1145/3603269.3604869, )  HetuMoE (nie2022hetumoe, )  DeepSpeed-MoE (DeepSpeedmoe, )  DeepSpeed-TED (DeepSpeedted, )  Lina (288705, )  ExFlow (yao2024exploiting, )  TA-MoE (chen2022ta, )  Aurora (li2024optimizing, )  LocMoE (li2024locmoe, )  Parm (10621327, ), ttnode] ] [Task Scheduling, wnode, l sep=5mm,calign=child edge, [ScMoE (cai2024shortcut, )  HiDup (HiDup, )  ScheMoE (10.1145/3627703.3650083, )  PipeMoE (10228874, )  EPS-MoE (qian2024eps, ), ttnode] ] ] [Expert Offloading, xnode, l sep=5mm, [Expert Prefetching, wnode, l sep=5mm,calign=child edge, [HOBBIT (tang2024hobbit, )  ExpertFlow (he2024expertflow, )  AdapMoE (zhong2024adapmoe, )  SiDA (du2024sida, )  MoE-Infinity (xue2024moe, )  Pre-gated MoE (hwang2024pre, )  MoE-Offloading (eliseev2023fast, )  EdgeMoE (yi2023edgemoe, ) DyNN-Offload (dynnoffload, )  Read-ME (readmoe, )  ProMoE(song2024promoe, ), ttnode] ] [Expert Caching, wnode, l sep=5mm,calign=child edge, [HOBBIT (tang2024hobbit, )  ExpertFlow (he2024expertflow, )  AdapMoE (zhong2024adapmoe, )  MoE-Infinity (xue2024moe, )  Fiddler (kamahori2024fiddler, )  EIO-MoE (yuan2024efficient, )  Pre-gated MoE (hwang2024pre, )  MoE-Offloading (eliseev2023fast, )  MoE-Deploy (huang2023towards, )  EdgeMoE (yi2023edgemoe, )  SwapMoE (kong2023serving, ), ttnode] ] [Expert Loading, wnode, l sep=5mm,calign=child edge, [HOBBIT (tang2024hobbit, )  AdapMoE (zhong2024adapmoe, )  EdgeMoE (yi2023edgemoe, ), ttnode] ] [CPU Assisting, wnode, l sep=5mm,calign=child edge, [HOBBIT (tang2024hobbit, )  Fiddler (kamahori2024fiddler, )  MoE-Lightning (cao2024moelight, ), ttnode] ] ] ]
@@
Due to the unique structure of MoE, many studies focus on accelerating inference at the system level by leveraging the sparse activation patterns inherent to this architecture. Typically, MoE models are deployed in two scenarios: cloud environments with multiple servers and edge environments with a single device. In cloud clusters, MoE models are distributed across many devices to enable parallel execution. In addition to traditional parallelization techniques such as data parallelism, tensor parallelism, and pipeline parallelism (narayanan2021efficient, ; rajbhandari2020zero, ; jia2019beyond, ), expert parallelism is a specialized method tailored specifically for MoE models. On edge devices, limited GPU memory often cannot accommodate all parameters of an MoE model, necessitating the offloading of some parameters to CPU memory or SSD storage. To address this, the expert-offloading technique has been developed to fully utilize the sparse activation patterns of experts for efficient execution. Figure 6 shows the detailed structure of this section.

### v2

@@ 3.2.2. Expert Quantization
Method Quantization Type Memory Reduction Accuracy Drop Inference Speedup Quantization Bits
Weight Activation
MC-MoE (li2023merge, ) 4.27x 3.8% 1.80x 1, 2, 3
MoE-CSP (kim2022says, ) 4.00x - 26.00x 4, 8
MoQE (kim2023mixture, ) 4.90x 0.97% - 2, 3, 4
QMoE (frantar2023qmoe, ) 20x 6.7% 0.95x 1, 2
CMoE (yuancompressed, ) 150x 23.81% - 1, 2, 4
MoE-MPTQS (imani2024mixture, ) - 0 $\sim$ 4.98% 1.00x $\sim$ 20.63x 4, 8
HOBBIT (tang2024hobbit, ) - 1% 1.35x 2, 4
EdgeMoE (yi2023edgemoe, ) 1.05x $\sim$ 1.18x 5% 1.11x $\sim$ 2.78x 2, 4, 8
Table 3. A comparison of quantization methods.
@@
In addition to model pruning, quantization is an effective technique for reducing model size by converting high-precision weights into low-precision. Given the redundancy of experts in MoE models, current research primarily focuses on quantizing the weights of the experts in MoE. As illustrated in Figure 4-(b), experts are quantized into appropriate low-precision versions using various methods. MC-MoE (huang2024mc, ) leverages the access frequency and activation weight to assess the importance of each expert. Along with the associated quantization loss, these metrics are used to determine the optimal quantization configuration for experts via the Integer Programming model. Specifically, the $i$-th expert’s access frequency is defined as $\phi_{i}=\frac{n_{i}}{N}$, and the activation weight is defined as $w_{i}=\frac{\sum_{j=1}^{N}\sigma_{j}}{N}$, where $n_{i}$ is the usage frequency, $\sigma_{j}$ is the routing weight, and $N$ represents the size of the calibration dataset. The quantization loss, $\epsilon_{ij}$ (computed using the Frobenius norm), is then determined for quantizing expert $i$ to $j$ bits ($j\in{1,2,3}$). Using these metrics, MC-MoE defines the objective function as $\sum_{i=1}^{n}\sum_{j=1}^{3}\phi_{i}^{\alpha}\cdot w_{i}^{\beta}\cdot(\epsilon_{ij}\cdot x_{ij})^{\gamma}$, which is minimized using Integer Programming to determine the optimal bit-width for each expert. MoE-CSP (kim2022says, ) quantizes expert weights to either 4 or 8 bits to reduce memory consumption in MoE models. Additionally, it designs specific CUDA kernels that handle the 4-bit/8-bit quantized weights and perform floating-point calculations to accelerate computations. MoQE (kim2023mixture, ) quantizes expert weights to 2 bits to address the memory and latency challenges of MoE models based on its observations that quantizing expert FFN layers to 2 bits does not significantly affect model quality, while quantizing other components, like self-attention, significantly hurts performance. Further advancements include QMoE (frantar2023qmoe, ) and CMoE (yuancompressed, ), which aggressively compress MoE model into just 1 bit. QMoE implements a highly scalable compression algorithm for large models and introduces a custom compression format along with bespoke GPU kernels for efficient on-the-fly computation. On the other hand, CMoE uses binary-weight networks to quantize model weights to 1 bit and applies learned step-size quantization to activations, quantizing them to 4 bits for MoE-based ASR models, enabling deployment on embedded devices. Moreover, some MoE-optimized systems leverage quantization for better system efficiency. MoE-MPTQS (imani2024mixture, ) and HOBBIT (tang2024hobbit, ) dynamically select quantized experts to replace original experts based on the current inputs, thereby reducing the expert loading cost on memory-limited devices. EdgeMoE (yi2023edgemoe, ) statistically determines the appropriate expert bit-width by profiling expert importance on a calibration dataset. Furthermore, QMoE-Benchmark (li2024examining, ) provides a benchmark for exploring various MoE structure-aware quantization heuristics, from coarse to fine granularity. The study reveals that different MoE structures (e.g., blocks, experts, linear layers) require different numbers of weight bits for effective and efficient quantization.
We summarize the main results reported by the methods discussed above in Table 3. From the table, we observe that quantization primarily benefits memory reduction, with most methods achieving more than a 4x reduction in memory usage. Some methods also lead to actual inference speedup, while others do not. For instance, QMoE even incurs a 5% overhead due to the absence of a dedicated 1-bit inference CUDA kernel implementation and hardware support. Furthermore, quantization typically causes some accuracy loss, with lower bit widths resulting in greater accuracy degradation. Therefore, when applying quantization to a model, it is crucial to strike a balance between memory consumption, accuracy, and inference speedup, taking into account the specific requirements and available hardware resources.
@@ 3.2.3. Expert Distillation
Knowledge distillation is another effective method for creating smaller, yet powerful models from larger ones. As shown in Figure 4-(c), knowledge distillation presents a promising solution to generate a compact, high-performance model from the original MoE model. LLaVA-MoD (shu2024llava, ) combines the MoE structure with knowledge distillation to efficiently train small multimodal large language models (s-MLLMs) from large ones (l-MLLMs). It first incorporates the MoE structure into the s-MLLM to balance computational efficiency and model performance. Then, it introduces two consecutive distillation stages, mimic distillation and preference distillation, to train the s-MLLM. Mimic distillation minimizes the Kullback-Leibler (KL) divergence between the output distributions of the s-MLLM and l-MLLM, enabling the s-MLLM to emulate the l-MLLM’s understanding. Preference distillation further refines the s-MLLM using Preference Optimization with additional datasets. DeepSpeed-MoE (DeepSpeedmoe, ) employs staged knowledge distillation to create a distilled version of its proposed PR-MoE model, called MoS. This method reduces model size while maintaining performance. Additionally, some studies focus on transferring the power of sparse models to dense models through knowledge distillation for more efficient deployment. For example, OneS (xue2022one, ) generates a dense model from a MoE model in two steps: knowledge gathering and knowledge distillation. Knowledge gathering merges experts into a single expert using four aggregation methods, including summation, averaging, top-k knowledge gathering, and Singular Value Decomposition (SVD) knowledge gathering. In the second step, knowledge distillation distills the merged model using the original MoE model. What’s more, MoE-KD (salinas2022knowledge, ) and CMoE (yuancompressed, ) distill MoE-based speech recognition models into dense models, accelerating the speech recognition process. Specifically, MoE-KD initializes the FFN module of the student dense model with the most frequently used expert from the teacher MoE model, then trains the FFN modules of the student model through layer-wise distillation. CMoE distills a binary dense model from the original model using quantization techniques. Switch Transformers (fedus2022switch, ) and ELSM (artetxe2021efficient, ) also explore distillation techniques to convert their sparse models into dense models. Additionally, to simplify the construction of MoE models, LaDiMo (kim2024ladimo, ) converts a dense model into a sparse MoE model via layer-wise distillation, where each expert learns to approximate the results of the original dense layers.
@@ 3.2.4. Expert Decomposition
As shown in Figure 4-(d), low-rank decomposition is an effective method for reducing model size by decomposing a large weight matrix into smaller matrices. MPOE (gao2022parameter, ) employs the matrix product operator (MPO), a tensor decomposition technique derived from quantum many-body physics, to decompose the expert weight matrix into a central tensor and several auxiliary tensors. The central tensor retains most of the parameters and core information of the original weight matrix, while the auxiliary tensors are much smaller and serve as complements to the central tensor. After decomposition, all experts in a layer share the same central tensor, thereby significantly reducing the total number of parameters in that layer. MC-SMoE (li2023merge, ) first groups the experts into several clusters and then merges each group into a single expert using a weighted sum. Low-rank decomposition is then applied to the merged experts to further reduce the model size. This approach is based on the observation that the merged experts have a lower rank compared to the original experts, making them more suitable for decomposition. MoE-$I^{2}$ (yang2024moe, ) identifies the importance of each expert and assigns higher ranks to more important experts while assigning lower ranks to less important ones for low-rank decomposition. The importance, $I_{i,j}$, of the $j$-th expert in the $i$-th layer, $e_{i,j}$, is determined by removing $e_{i,j}$ and calculating the performance loss compared to the original model. The SVD rank, $r_{i,j}$, for $e_{i,j}$ is then computed as $r_{i,j}=\left\lfloor\frac{(I_{i,j}+\epsilon)^{\alpha}}{\sum_{j=1}^{M_{i}}(I_{i,j}+\epsilon)^{\alpha}}\right\rfloor\cdot R_{a}\cdot M_{i}$, where $\epsilon$ and $\alpha$ are hyperparameters, $R_{a}$ is the compression ratio, and $M_{i}$ is the number of experts in layer $i$.
3.3. Algorithm Improvement
In this part, we conduct an in-depth exploration of two other strategies for improving the MoE model inference algorithm.
Figure 5. Algorithm improvement on expert layers.
@@ 3.3.1. Dynamic Gating
Given the significant progress made in utilizing the sparsity of MoE models, specific strategies can be employed to further exploit this sparsity in the inference process. Due to the vertical parallel structure of MoE experts, dynamically reducing the number of experts activated for each token clearly presents an effective strategy. Figure 5-(a) illustrates the MoE layer’s calculation process after the skip mechanism is applied, in contrast to the original process.
Li et al. (li2023adaptive, ) proposes a self-adaptive gating mechanism that dynamically determines the number of experts required for each token, based on the expert probability distribution. This method enhances training efficiency while preserving the sparsity of the MoE model and further reduces training time through curriculum learning. DynMoE (guo2024dynamic, ) introduces two innovative methods for expert activation: a top-down gating approach that enables flexible per-token expert allocation, and an adaptive mechanism that dynamically determines the number of experts needed for each token based on computational requirements. XMoE (yang2024xmoe, ) employs a strategy of using smaller, but more experts, with dynamic expert activation based on threshold values to balance computational efficiency and model performance. AdapMoE (zhong2024adapmoe, ) is a co-design framework aimed at improving inference efficiency on edge devices. It adaptively applies three techniques: expert activation, expert prefetching, and GPU cache allocation. The comprehensive comparison of performance and methods of these works is presented in Table 4. DA-MoE (aghdam2024damoedynamicexpertallocation, ) introduces a token importance prediction method derived from attention mechanisms to guide expert allocation. By assigning experts based on token importance scores, their approach achieves strong performance on the GLUE benchmark while demonstrating effective scaling with increased expert count.
Method FLOPs Reduction Speedup Threshold Strategy Load Balance PR.
Fixed top-k gating 0% 1.0x
Li et al.(li2023adaptive, ) 38.2% 1.32x accumulative probability soft on top-1
DynMoE(guo2024dynamic, ) 9% 1.37x single expert probability
XMoE(yang2024xmoe, ) 75% - accumulative probability
AdapMoE(zhong2024adapmoe, ) 25% of experts 1.35x performance perturbation
Table 4. A comparison of dynamic gating methods. PR. indicates performance Retention. Specially, Li et al. (li2023adaptive, ) only uses soft load balance constraints on top-1 gating. No specific acceleration ratio was provided in XMoE (yang2024xmoe, ).
@@ 3.3.2. Sparse to Dense
In certain scenarios, dense models offer unique advantages due to their smaller number of parameters. Therefore, transforming MoE models into dense target models all at once can achieve maximum sparsity while maintaining model performance. Most approaches use knowledge distillation to achieve this sparse-to-dense conversion, as illustrated in Figure 5-(b).
XFT (ding2024xft, ) proposes a novel method to supervise fine-tune dense LLMs. They first generate a sparse-upcycled MoE model, and then transform it back into an efficient dense LLM of the same size and structure through a learnable merging mechanism. Switch Transformers (fedus2022switch, ) explores distillation techniques to convert a sparse model into a dense one, demonstrating that the dense model can retain over 30% of its performance even after compressing 97% of the parameters from the sparse MoE model. ELSM (Artetxe et al., 2021) demonstrates that dense student models distilled from sparse MoE teachers can match and even surpass the teacher’s performance. OneS (xue2022one, ) employs four distinct methods to generate the dense model, including summation, averaging, top-k Knowledge Gathering, and Singular Value Decomposition Knowledge Gathering. The model is then refined to reduce noise by gathering sparse knowledge. TSEP (chen2022task, ) progressively eliminates non-expert components based on specific downstream tasks, ultimately converting the sparse MoE model into a dense counterpart. EWA (huang2023experts, ) replaces FFNs with specially designed MoEs during training before reverting to dense ViT for inference. AdaMoLE (liu2024adamolefinetuninglargelanguage, ) combining the Low-Rank Adaptation (LoRA) structure, a dedicated network is used to adjust the activation threshold for different task complexities. It has shown superior performance to baseline in many natural language processing tasks, especially in some commonsense reasoning tasks.
@@ 4. System-level Optimizations
Communication, wnode, l sep=5mm,calign=child edge, [Tutel (hwang2023tutel, )  Janus (10.1145/3603269.3604869, )  HetuMoE (nie2022hetumoe, )  DeepSpeed-MoE (DeepSpeedmoe, )  DeepSpeed-TED (DeepSpeedted, )  Lina (288705, )  ExFlow (yao2024exploiting, )  TA-MoE (chen2022ta, )  Aurora (li2024optimizing, )  LocMoE (li2024locmoe, )  Parm (10621327, ), ttnode] ] [Task Scheduling, wnode, l sep=5mm,calign=child edge, [ScMoE (cai2024shortcut, )  HiDup (HiDup, )  MoESys (10528887, )  ScheMoE (10.1145/3627703.3650083, )  PipeMoE (10228874, )  EPS-MoE (qian2024eps, ), ttnode] ] ] [Expert Offloading, xnode, l sep=5mm, [Expert Prefetching, wnode, l sep=5mm,calign=child edge, [HOBBIT (tang2024hobbit, )  ExpertFlow (he2024expertflow, )  AdapMoE (zhong2024adapmoe, )  SiDA (du2024sida, )  MoE-Infinity (xue2024moe, )  Pre-gated MoE (hwang2024pre, )  Mixtral-Offloading (eliseev2023fast, )  EdgeMoE (yi2023edgemoe, ) DyNN-Offload (dynnoffload, )  Read-ME (readmoe, )  ProMoE(song2024promoe, ), ttnode] ] [Expert Caching, wnode, l sep=5mm,calign=child edge, [HOBBIT (tang2024hobbit, )  ExpertFlow (he2024expertflow, )  AdapMoE (zhong2024adapmoe, )  MoE-Infinity (xue2024moe, )  Fiddler (kamahori2024fiddler, )  EIO-MoE (yuan2024efficient, )  Pre-gated MoE (hwang2024pre, )  Mixtral-Offloading (eliseev2023fast, )  MoE-Deploy (huang2023towards, )  EdgeMoE (yi2023edgemoe, )  SwapMoE (kong2023serving, )  CacheMoE (skliar2024mixture, ), ttnode] ] [Expert Loading, wnode, l sep=5mm,calign=child edge, [HOBBIT (tang2024hobbit, )  AdapMoE (zhong2024adapmoe, )  EdgeMoE (yi2023edgemoe, ), ttnode] ] [CPU Assisting, wnode, l sep=5mm,calign=child edge, [HOBBIT (tang2024hobbit, )  Fiddler (kamahori2024fiddler, )  MoE-Lightning (cao2024moelight, ), ttnode] ] ] ]
@@
Due to the unique structure of MoE, many studies focus on accelerating inference at the system level by leveraging the sparse activation patterns inherent to this architecture. Typically, MoE models are deployed in two scenarios: cloud environments with multiple powerful servers and edge environments with a single device. In cloud clusters, MoE models are distributed across many devices to enable parallel execution. In addition to traditional parallelization techniques such as data parallelism, tensor parallelism, and pipeline parallelism (narayanan2021efficient, ; rajbhandari2020zero, ; jia2019beyond, ), expert parallelism is a specialized method tailored specifically for MoE models. On edge devices, limited GPU memory often cannot accommodate all parameters of an MoE model, necessitating the offloading of some parameters to CPU memory or SSD storage. To address this, the expert-offloading technique has been developed to fully utilize the sparse activation patterns of experts for efficient execution. Figure 6 shows the detailed structure of this section.